The consumption expenditure in the economy is $200mn. The investment expenditure in the economy is $100 mn. The Government spending is $250 mn. The exports are $100mn and imports are $300mn. What is the GDP of the country?
Answer #1

Answer

GDP is found by the following equation

Y=C+I+G+NX

Y=GDP

C=consumption

G=govenment spending

I=investment

NX=net export =export - import=100-300=-200

Y=200+100+250+(-200)

Y=350

The GDP is $350